Kinetic energy of a particle moving in a straight line varies with time t as k=4t•2. The force acting on the particle

Grade:12th pass

2 Answers

Vikas TU
7 years ago
Force is also defined as:
That is:
F = dP/dt
and P^2/2m = K = 4t^2
we get,
 P = t*root(8m)
F = dP/dt => root(8*m) N.
venkata ramana
7 years ago
sir i don’t understand plese send me formula once more.................................................................